# HLPD: Aligning LLMs to Human Language Preference for Machine-Revised Text Detection

### v2

@@ 2 Methodology
We present a novel method, Human Language Preference Detection (HLPD), for detecting both machine‑revised and fully generated texts by aligning the scoring model to human writing styles. This alignment improves robustness in the black-box setting, where the specific LLM is unknown, and improves sensitivity to outputs from state-of-the-art LLMs in diverse revision and generation scenarios. For more details of those scenarios, please refer to Section 3.1 and Appendix C.
@@ 3.1 Experiment Settings
Constructing the Single-Task Dataset. Following ImBD’s methodology for revision tasks and generation task, datasets were constructed based on three distinct tasks: Rewriting, Expand, Polish and generate. The process involved a two-stage pipeline: (1) Revision instructions generation (2) Paragraph revision under generated instructions. See Appendix C.1, C.2 for details of all tasks.
Constructing the Adversarial Multi-Task Dataset. Adversarial Dataset Construction. To evaluate model robustness under realistic conditions, we constructed a challenging adversarial dataset inspired by Zhao et al. (2024). We developed a five-dimensional prompt generator that creates complex instructions by combining: (1) a core revision goal, (2) a target style, (3) instructions to add human-like qualities for detection evasion, (4) operational constraints, and (5) auxiliary user requests. Our generation pipeline consists of two stages: first, these five components are used to prompt DeepSeek-R1, which generates a consolidated instruction. Second, this instruction, along with a human-written text, is fed to a target LLM to produce the final adversarial sample. This methodology expanded our prompt pool to 750 unique prompts. The detailed pipeline is illustrated in Appendix C.3.
@@
To ensure the robustness of our findings, all key experiments were repeated five times with different random seeds.Figure 3 presents the ROC curves for detection performance. The notation HLPD†refers to a baseline version implemented without any optimizations. Further implementation and traing details are provided in Appendix B.
@@ 3.2 Effectiveness & Efficiency (RQ1)
Detection Performance for Single-task Revision. As shown in Table 1, in ImBD’s rewrite task, our method outperforms ImBD by 15.11% relative improvement in detecting GPT-series outputs. Compared to the ReMoDetect and Ghostbuster, our method achieves average improvements of 9.27% and 3.21% , respectively. Additionally, our method achieves a 10.57% relative improvement over GPTZero. This indicates that our approach is highly training-efficient, attaining superior performance with a limited amount of data compared to models trained on significantly large datasets. See Appendix E Table 1 for the performance on polish task. Furthermore, our method gets the highest score on four open-source models. See Appendix E Table 2, 5 and 6 for detailed results.
@@
Multilingual Detection Results. To evaluate the performance of HLPD in multilingual text scenarios, we generated Chinese, Portuguese, and Spanish dataset variants. As shown in Appendix E Figure 5, HLPD exhibits superior detection capabilities: achieving a 10% relative advantage over ImBD and gaining an average improvement of 33% against RoBERTa-large across all language versions.
Inference Efficiency. As shown in Table 1, HLPD achieves an average inference time of 0.72 seconds per 1,000 words, matching the efficiency of Fast-DetectGPT while delivering superior detection performance on an L20.
@@ 3.3 Robustness Studies(RQ2)
Robustness in Diverse Revision Goals. HLPD’s robustness against diverse revision goals is evident in Appendix E Table 3. Evaluated on the XSum dataset revised by six LLMs, HLPD achieved the highest average AUROC across all three revision tasks. This performance represents a significant 32.38% improvement over Fast-DetectGPT and a 4.11% relative gain over ImBD, demonstrating HLPD’s strong capability in multi-task revision and multi-context detection. For further details, see Appendix E Table 7.
Robustness in Multiple Revision Iterations. To assess HLPD’s robustness to multiple revision iterations, we evaluated its performance on texts undergoing one to five revisions from two LLMs: GPT-4o and Gemini-2.5-Pro. As shown in Appendix Table 4, HLPD consistently maintained highest average AUROC scores, indicating HLPD’s detection efficacy even as texts undergo extensive iterative machine revisions.
@@ 3.5 Humanization Adaptive Attack Experiments
The selected perturbation $\tilde{x}_{m}$ can be applied to additional iterations for further reductions in detection probability. However, it is important to note that repeating this process might introduce more substantial semantic deviations from the original text. Please see Appendix D for the process flow diagram.
Experiment Settings. We generate our text samples via a two-stage process. First, to create the base machine-revised text, we use GPT-3.5-Turbo to polish 100 paragraphs sourced from the XSum dataset. Second, for each revised paragraph, we employ a perturbation-and-selection method: T3-B5 generates 100 candidate perturbations, and our scoring model selects the one with the highest preference score as the final "humanized" output for each humanization.
Effectiveness Against Commercial Detectors. As shown in Appendix D Figure 4, our method reduces the detection AUROC by an average of 5% per iteration, culminating in a 20% total reduction after four iterations compared to the original machine-revised text. On the commercial GPTZero platform, this corresponds to a significant 74% decrease in the predicted AI probability. These results highlight the efficacy of our iterative humanization technique in evading commercial detection systems.
Effectiveness Against Open-Source Detectors. The method demonstrates similar effectiveness against open-source detectors. On the Imbue¹ demo, our 4-iteration humanized text reduced the AI probability score by 44.5% (from 81.8% to 37.3%). Similarly, on the Ghostbuster demo, the probability dropped by 61% (from 87% to 26%). See detailed examples of the original and humanized texts in Appendix D.3.
@@ B.3 Multi-Task Adversarial Revised text generation
Figure 6: Humanization of Machine-Revised Texts. We generate perturbed versions of the original machine-revised text and then use the HLPD model to select, fool the online detection platform GPTZero, making it think it was written by a human.
@@ C.2 Additional Humanizing results
Humanizing against GPTZero. As illustrated in Figure  7, our iterative humanization technique proves highly effective at evading AI detection. Each iteration reduces the detection AUROC by an average of 5%, culminating in a 20% total reduction after four cycles compared to the original machine-revised text. This corresponds to a remarkable 74% decrease in the AI probability score assigned by the commercial GPTZero platform, underscoring the practical efficacy of our method against contemporary detection systems.
@@ C.3 Example of humanized Machine-revised text
Detailed Results for Machine-Polished Text Detection. As presented in Table 1, 2 our method demonstrates superior performance in detecting texts revised by GPT-3.5-turbo and gpt-4o-2024-05-13. It consistently achieved the highest accuracy across all evaluated datasets—XSUM, SQuAD, PubMed, Wikitext and WritingPrompts—outperforming all competing methods.
@@
Robustness to Diverse Revision Goals. As Detailed in Table 2, 3, HLPD shows strong robustness against diverse revision goals. On the XSum dataset, revised by six different LLMs across three revision tasks, HLPD achieved the highest average AUROC. This represents a 32.38% improvement over Fast-DetectGPT and a 4.11% gain over ImBD, underscoring its strong multi-task detection capabilities.
@@
Detailed results on open-source model revised text. The results presented in Tables 5 and 6 underscore the clear superiority of our method for detecting content from open-source LLMs. This superior performance holds true regardless of the generative model, proving effective against Qwen2-7B, Llama-3-8B, Mistral-7B, and Deepseek-7B.

### v3

@@ 2 Methodology
We present a novel method, Human Language Preference Detection (HLPD), for detecting both machine‑revised and fully generated texts by aligning the scoring model to human writing styles. This alignment improves robustness in the black-box setting, where the specific LLM is unknown, and improves sensitivity to outputs from state-of-the-art LLMs in diverse revision and generation scenarios. For more details of those scenarios, please refer to Section 3.1 and Appendix B.
@@ 3.1 Experiment Settings
Constructing the Single-Task Dataset. Following ImBD’s methodology for revision tasks and generation task, datasets were constructed based on four distinct tasks: Rewriting, Expand, Polish and Generate. The process involved a two-stage pipeline: (1) Revision instructions generation (2) Paragraph revision under generated instructions. See Appendix B.1, B.2 for details of all tasks.
Constructing the Adversarial Multi-Task Dataset. To evaluate model robustness under realistic conditions, we constructed a challenging adversarial dataset inspired by Zhao et al. (2024). We developed a five-dimensional prompt generator that creates complex instructions by combining: (1) a core revision goal, (2) a target style, (3) instructions to add human-like qualities for detection evasion, (4) operational constraints, and (5) auxiliary user requests. Our generation pipeline consists of two stages: first, these five components are used to prompt DeepSeek-R1, which generates a consolidated instruction. Second, this instruction, along with a human-written text, is fed to a target LLM to produce the final adversarial sample. This methodology expanded our prompt pool to 750 unique prompts. The detailed pipeline is illustrated in Appendix B.3.
@@
[…]
@@ 3.2 Effectiveness & Efficiency (RQ1)
Detection Performance for Single-task Revision. As shown in Table 1, in ImBD’s rewrite task, our method outperforms ImBD by 15.11% relative improvement in detecting GPT-series outputs. Compared to the ReMoDetect and Ghostbuster, our method achieves average improvements of 9.27% and 3.21%, respectively. Additionally, our method achieves a 10.57% relative improvement over GPTZero. This indicates that our approach is highly training-efficient, attaining superior performance with a limited amount of data compared to models trained on significantly large datasets. See Appendix D Table 5 for the performance on polish task. Furthermore, our method gets the highest score on four open-source models. See Appendix D Table 6, 9, and 10 for detailed results.
@@
Multilingual Detection Results. To evaluate the performance of HLPD in multilingual text scenarios, we generated Chinese, Portuguese, and Spanish dataset variants. As shown in Appendix D Figure 8, HLPD exhibits superior detection capabilities: achieving a 10% relative advantage over ImBD and gaining an average improvement of 33% against RoBERTa-large across all language versions.
Inference Efficiency. As shown in Table 1, HLPD achieves an average inference time of 0.72 seconds per 1,000 words, matching the efficiency of Fast-DetectGPT while delivering superior detection performance on L20 GPU cards.
@@ 3.3 Robustness Studies(RQ2)
Robustness in Diverse Revision Goals. HLPD’s robustness against diverse revision goals is evident in Appendix D Table 7. Evaluated on the XSum dataset revised by six LLMs, HLPD achieved the highest average AUROC across all three revision tasks. This performance represents a significant 32.38% improvement over Fast-DetectGPT and a 4.11% relative gain over ImBD, demonstrating HLPD’s strong capability in multi-task revision and multi-context detection. For further details, see Appendix D Table 11.
Robustness in Multiple Revision Iterations. To assess HLPD’s robustness to multiple revision iterations, we evaluated its performance on texts undergoing one to five revisions from two LLMs: GPT-4o and Gemini-2.5-Pro. As shown in Appendix D Table 8, HLPD consistently maintained highest average AUROC scores, indicating HLPD’s detection efficacy even as texts undergo extensive iterative machine revisions.
@@ 3.5 Humanization Adaptive Attack Experiments
The selected perturbation $\tilde{x}_{m}$ can be applied to additional iterations for further reductions in detection probability. However, it is important to note that repeating this process might introduce more substantial semantic deviations from the original text. Please see Appendix C for the process flow diagram.
Experiment Settings. We generate our text samples via a two-stage process. First, to create the base machine-revised text, we use GPT-3.5-Turbo to polish 100 paragraphs sourced from the XSum dataset. Second, for each revised paragraph, we employ a perturbation-and-selection method: T3-B5 generates 100 candidate perturbations, and our scoring model selects the one with the highest preference score as the final “humanized” output for each humanization.
Effectiveness Against Commercial Detectors. As shown in Appendix C Figure 7, our method reduces the detection AUROC by an average of 5% per iteration, culminating in a 20% total reduction after four iterations compared to the original machine-revised text. On the commercial GPTZero platform, this corresponds to a significant 74% decrease in the predicted AI probability. These results highlight the efficacy of our iterative humanization technique in evading commercial detection systems.
Effectiveness Against Open-Source Detectors. The method demonstrates similar effectiveness against open-source detectors. On the ImBD demo, our 4-iteration humanized text reduced the AI probability score by 44.5% (from 81.8% to 37.3%). Similarly, on the Ghostbuster demo, the probability dropped by 61% (from 87% to 26%). See detailed examples of the original and humanized texts in Appendix C.3.
@@ B.3 Multi-Task Adversarial Revised text generation
We generate perturbed versions of the original machine-revised text and then use the HLPD model to select, fool the online detection platform GPTZero, making it think it was written by a human.
Figure 6: Humanization of Machine-Revised Texts.
@@ C.2 Additional Humanizing results
Humanizing against GPTZero. As illustrated in Figure 7, our iterative humanization technique proves highly effective at evading AI detection. Each iteration reduces the detection AUROC by an average of 5%, culminating in a 20% total reduction after four cycles compared to the original machine-revised text. This corresponds to a remarkable 74% decrease in the AI probability score assigned by the commercial GPTZero platform, underscoring the practical efficacy of our method against contemporary detection systems.
@@ C.3 Example of humanized Machine-revised text
Detailed Results for Machine-Polished Text Detection. As presented in Table 5 and 6 our method demonstrates superior performance in detecting texts revised by GPT-3.5-turbo and gpt-4o-2024-05-13. It consistently achieved the highest accuracy across all evaluated datasets—XSUM, SQuAD, PubMed, Wikitext and WritingPrompts—outperforming all competing methods.
@@
Robustness to Diverse Revision Goals. As detailed in Table 6, 7, HLPD shows strong robustness against diverse revision goals. On the XSum dataset, revised by six different LLMs across three revision tasks, HLPD achieved the highest average AUROC. This represents a 32.38% improvement over Fast-DetectGPT and a 4.11% gain over ImBD, underscoring its strong multi-task detection capabilities.
@@
Detailed results on open-source model revised text. The results presented in Tables 9 and 10 underscore the clear superiority of our method for detecting content from open-source LLMs. This superior performance holds true regardless of the generative model, proving effective against Qwen2-7B, Llama-3-8B, Mistral-7B, and Deepseek-7B.